A researcher wishes to see whether the mean of the heart rates (in beats per minutes) of smokers are different from the means of heart rates of people who do not smoke. The two samples are selected and the data are shown below. Using alfa = 0.05, is there enough evidence to support the claim?

 

 

Smoker

Non-smoker

n

60

60

Mean

79

86

s

32

38


1
Expert's answer
2022-01-10T13:34:52-0500

"H_0:\mu_1=\mu_2" , the mean of the heart rates of smokers is not different from the means of heart rates of people who do not smoke

"H_a:\mu_1\neq\mu_2" , the mean of the heart rates of smokers is not different from the means of heart rates of people who do not smoke


"t=\frac{\mu_1-\mu_2}{\sqrt{s_1^2/n_1+s^2_2/n_2}}=-1.091"


"df=\frac{(s_1^2/n_1+s^2_2/n_2)^2}{\frac{s_1^4}{n_1^2(n_1-1)}+\frac{s_2^4}{n_262(n_2-1)}}=114.68"


critical value:

"t_{crit}=1.981"


Since "|t|<t_{crit}" we accept null hypothesis. The mean of the heart rates of smokers is not different from the means of heart rates of people who do not smoke.